# Attention Layers Add Into Low-Dimensional Residual Subspaces

### v2

@@ Abstract
Transformer architectures, and their attention mechanisms in particular, form the foundation of modern large language models. While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are confined to a surprisingly low-dimensional subspace, where about 60% of the directions account for 99% of the variance–a phenomenon that is consistently observed across diverse model families and datasets, and is induced by the attention output projection matrix. Critically, we find this low-rank structure as a key factor of the prevalent dead feature problem in sparse dictionary learning, where it creates a mismatch between randomly initialized features and the intrinsic geometry of the activation space. Building on this insight, we propose a subspace-constrained training method for sparse autoencoders (SAEs), initializing feature directions into the active subspace of activations. Our approach reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features, and can further extend to other sparse dictionary learning methods. Our findings provide both new insights into the geometry of attention and practical tools for improving sparse dictionary learning in large language models.
@@ 1 Introduction
However, scaling these approaches presents practical difficulties (Templeton et al., 2024; Gao et al., 2024; Mudide et al., 2025). As models and feature dictionaries grow, the number of parameters increases rapidly, driving up computational costs. At the same time, the prevalence of dead features leads to substantial waste in computation and memory (Templeton et al., 2024; Kissane et al., 2024), limiting the efficiency of interpretability methods. In this work, we identify low-rank activation structure as a major driver of dead features (Section 5.1).
In Section 4, we show that attention outputs exhibit a remarkably strong low-rank structure compared to multilayer perceptron (MLP) outputs and residual streams. Through singular value decomposition and intrinsic dimensionality analyses (Guth et al., 2023; Staats et al., 2025), we demonstrate that this phenomenon holds universally across layers, datasets, and model families—Llama 3.1 (Dubey et al., 2024), Gemma 2 (Rivière et al., 2024), and Qwen 3 (Yang et al., 2025), which is consistent with the universality hypothesis (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025). We further trace the origin of this low-rank structure to the anisotropy of the output projection matrix $W^{O}$, which compresses the multi-head outputs into a lower-dimensional subspace.
In Section 5, we investigate how the low-rank nature of attention outputs interacts with SAE training. By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low intrinsic dimensionality strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations. Drawing inspiration from Phan et al. (2025)’s principal component initialization for the first network layer, we propose Active Subspace Initialization, which aligns SAE features with the active subspace of activations, substantially reducing dead features while improving reconstruction. Following Lindsey et al. (2024a) and Gao et al. (2024), we conduct scaling experiments, which further reveal that ASI achieves superior reconstruction across feature counts, and when combined with SparseAdam [^1] https://docs.pytorch.org/docs/stable/generated/torch.optim.SparseAdam.html, it achieves the best reconstruction in large scale and reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features trained on Llama-3.1-8B (Dubey et al., 2024).
@@ 2 Related Work
2.1 Low-Rankness in Attention Mechanisms
Prior work has investigated various notions of “low-rankness” within attention mechanisms.: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach & Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
@@ 2.1 Low-Rankness in Attention Mechanisms
2.2 Linear Representation Hypothesis and Sparse Dictionary Learning Methods
The linear representation hypothesis suggests that high-level concepts align with linear directions in representation space (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024). Motivated by this view, various sparse dictionary learning methods have been developed for interpretability, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025). These methods decompose activations into sparse feature combinations, while differing in their strategies for predicting or approximating feature activations. The hypothesis has been validated across diverse model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024).
@@ 2.4 PCA-Inspired Network Initialization
A common practice applies PCA to input data for dimensionality reduction before network training (Hastie et al., 2009; Montavon et al., 2012; Jolliffe, 1986; Bishop & Nasrabadi, 2007). Recently, Phan et al. (2025) proposed PCsInit, which initializes the first layer weights of networks with top principal components of data—embedding the PCA transform directly into the network. This provides the model with a superior parameter set (Gu et al., 2025), boosting performance by construction.
@@ 3.1 Multi-Head Self-Attention and Notations
$Q_{i}=XW^{Q}_{i},\quad K_{i}=XW^{K}_{i},\quad V_{i}=XW^{V}_{i},\quad W^{Q}_{i},W^{K}_{i},W^{V}_{i}\in\mathbb{R}^{d\times d_{h}},$
@@
$A_{i}=\mathrm{softmax}\left(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{h}}}\right),\quad Z_{i}=A_{i}V_{i}.$
Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated output of all attention heads (Nanda & Bloom, 2022). The final attention output is computed by applying the output projection:
$O=ZW^{O},\quad W^{O}\in\mathbb{R}^{d\times d}.$
This formulation shows that $O$ can be viewed as the sum of low-dimensional outputs from each head, projected into the residual stream space. $O$ is the attention block’s contribution to the residual stream.
@@ 3.2 TopK Sparse Autoencoders
$\hat{x}=W_{d}z+b_{d}.$
The model is trained to minimize the reconstruction loss, optionally augmented with an auxiliary penalty to prevent dead latents:
$\mathcal{L}_{\text{TopK-SAE}}=\|x-\hat{x}\|_{2}^{2}+\alpha\cdot\mathcal{L}_{\text{aux}},$
@@ 4 Low-Rank Structure of Attention Outputs
Figure 1: Across layers, model families and datasets, attention outputs exhibit dramatically lower intrinsic dimensions (details in Section 4.1) than residual streams and MLP outputs, showing that the attention layer writing into a low dimensional subspace of residual stream is a universal phenomenon.
@@
We begin by presenting our central empirical finding: in Transformer models, attention outputs consistently display the strongest low-rank structure compared to MLP outputs and residual streams. As shown in Figure 1, attention outputs have a significantly lower intrinsic dimension. This phenomenon is remarkably robust, holding across different intermediate layers, model families and datasets. These observations highlight that the attention block modifies a subspace of the residual stream, while the MLP operates nearly on the full space.
4.1 Quantifying Low-Rankness with Relative Singular Values
@@
We consider activation matrix $\widetilde{A}\in\mathbb{R}^{n\times d}$, where each row corresponds to one token’s activation vector, and $n$ is the number of data points, while $d$ is the dimensionality of the activation space (e.g., the hidden size of the model). Unless otherwise specified, $\widetilde{A}$ refers to the mean-centered activations. We refer readers to Appendix D for more details of activations sources.
To quantify the rank of data, we perform singular value decomposition (SVD) on $\widetilde{A}$:
$\widetilde{A}=U\Sigma V^{\top},$
@@
To analyze the intrinsic dimension of these activations, we compute the smallest integer $k$ such that:
$\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{i=1}^{r}\sigma_{i}^{2}}\geq\tau,$
for a given threshold $\tau\in(0,1)$. Since SVD yields the optimal low-rank approximation in terms of reconstruction error, this $k$ provides a principled way to assess how concentrated the activations are in a low-dimensional subspace [^3] We use 0.99 as the threshold in main text, results of some other thresholds are provided in Appendix F, with no influence to the conclusion. (Figure 1). We further compute the fraction of delta downstream loss recovered by different number of components (Section 4.2). These metrics complement our central findings, offering a numerical characterization of low-rankness.
For each type of activations, we collect 10 million activation vectors [^4] In rare cases, outlier activations inflate variance along some directions, making variance-based dimensionality estimates unreliable; to address this, we filter activations with a norm exceeding $5\sigma$ from the mean.. We empirically verify that this sample size is sufficient to produce stable and reproducible singular spectrum analysis.
4.2 Empirical Evidence of Low-Rank Structure
Figure 2: Results for 1010M tokens from SlimPajama (Soboleva et al., 2023) fed into Llama-3.1-8B. (a) The attention outputs are the most low-rank, as indicated by the sharpest decay in singular values. (b) Fraction of loss recovered using varying numbers of top singular components.
We draw our findings from three lines of evidence:
Low Intrinsic Dimension of Attention Outputs
Attention outputs has an intrinsic dimension of around 60% of the total dimensionality. In contrast, MLP outputs and the residual streams show much higher intrinsic dimensions above 90% (Figure 1).
Rapid Singular Spectral Decay in Attention Outputs
This is quantitatively evidenced by the number of components retaining significant energy: only 74.7% singular values exceed 1% of the maximum in attention outputs, versus 100.0% for MLP outputs and residual streams (Figure 2(a)).
Efficient Downstream Loss Recovery
Compared to zero ablation, attention outputs requires only 39.1% of dimensions to recover over 99% of the downstream loss, versus 95.3% and 96.9% for MLP outputs and residual streams (Figure 2(b)).
More results of these metrics across different layers, models and datasets are shown in Appendix E.
4.3 Low-Rankness of Attention Outputs Results from the Output Projection Matrix
Among all activation types, attention outputs consistently exhibits the most rapid spectral decay. To investigate whether this low-rank structure originates from the attention heads ($Z$), the output projection matrix ($W^{O}$), or their interaction, we perform a decomposition-based analysis.
Recall that the attention output is computed as $O=ZW^{O}$, where $Z\in\mathbb{R}^{n\times d}$ is the concatenated output of all attention heads, and $W^{O}\in\mathbb{R}^{d\times d}$ is a learned linear projection. To understand how the variance in $O$ (singular value spectra of O) is shaped, we analyze the variance of $O$ along an arbitrary unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
@@ 4.3 Low-Rankness of Attention Outputs Results from the Output Projection Matrix
This expression highlights that the variance along $e$ is determined by two factors: the norm of $W^{O}e$ and the variance of $Z$ projected onto the direction $W^{O}\hat{e}$. Specifically, we can rewrite the variance as:
$\mathrm{Var}(O\hat{e})=\mathrm{Var}(Z\hat{v})\cdot\|v\|_{2}^{2}\quad\text{where}\quad v=W^{O}\hat{e},\ \hat{v}=\frac{v}{\|v\|_{2}}.$
We refer to $\mathrm{Var}(Zv)$ as the contribution of Z, capturing how much variance the head output $Z$ provides in that direction, and $\|v\|_{2}^{2}$ as the contribution of $W^{O}$, measuring how much the output projection $W^{O}$ scales or suppresses that direction.
We compute and visualize both quantities across a set of directions aligned with the right singular vectors of attention output, as shown in Figure 3. Our analysis reveals that the low-rank structure of attention outputs $O$ arises primarily from the anisotropy of $W^{O}$, which heavily compresses the output space into a lower-dimensional subspace. From a mechanistic perspective, an intuitive way to see this is that although each attention head contributes a $d_{\text{head}}$-dimensional subspace, the superposition of heads (Jermyn et al., 2024; He et al., 2025) inherently leads to overlaps among these subspaces. We note the output of the $i^{th}$ head as $\text{head}_{i}$. Consequently, the dimension of the MHSA output satisfies
@@
$\dim\!\left(\bigcup_{i}\text{span}(\text{head}_{i})\right)\;\leq\;\sum_{i}\dim\!\big(\text{span}(\text{head}_{i})\big)\;=\;d_{\text{head}}\cdot n_{\text{head}}\;(\;=\;d_{\text{model}}\ \text{in standard MHSA)}.$

### v3

# Dimensional Collapse in Transformer Attention Outputs: A Challenge for Sparse Dictionary Learning

@@ Abstract
Transformer architectures, and their attention mechanisms in particular, form the foundation of modern large language models. While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are in fact confined to a surprisingly low-dimensional subspace, with an effective dimensionality of only about $60\%$ of the full space. In contrast, MLP outputs and residual streams remain much closer to full-rank, exhibiting effective ranks around $90\%$. This striking dimensional discrepancy is consistently observed across diverse model families and datasets, and is strongly shaped by the attention output projection matrix. Critically, we find this low-rank structure as a key factor of the prevalent dead feature problem in sparse dictionary learning, where it creates a mismatch between randomly initialized features and the intrinsic geometry of the activation space. Building on this insight, we propose a subspace-constrained training method for sparse autoencoders (SAEs), initializing feature directions into the active subspace of activations. Our approach reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features, and can further extend to other sparse dictionary learning methods. Our findings provide both new insights into the geometry of attention and practical tools for improving sparse dictionary learning in large language models.
Machine Learning, ICML
Figure 1: (left) Attention outputs exhibit pronounced low-rank structure compared to residual streams and MLP outputs, indicating that the attention layer writes into a subspace of the residual stream. (right) Low effective dimensionality of activations is a root cause of dead features in sparse dictionary learning methods. Setting feature directions in the active subspace mitigates this issue.
@@ 1 Introduction
However, scaling these approaches presents substantial practical challenges (Templeton et al., 2024; Gao et al., 2024; Mudide et al., 2025). As models and feature dictionaries grow, the parameter count increases rapidly, leading to significant computational overhead. Moreover, a large fraction of learned features remain inactive, resulting in considerable waste in both computation and memory (Templeton et al., 2024; Kissane et al., 2024). In this work, we identify the low-rank structure of the activations as a primary driver of dead features (Section 5.1).
Through singular value decomposition and effective dimensionality analyses (Roy and Vetterli, 2007; Staats et al., 2025), we show for the first time that the outputs of multi-head self-attention in transformer-based language models exhibit a remarkably low-rank structure (Section 4). Compared to multilayer perceptron (MLP) outputs and residual streams, attention outputs consistently concentrate in a low-dimensional subspace. We show that this behavior is robust across layers, datasets, and model families, including GPT-2 (Radford et al., 2019), Llama 3.1 (Dubey et al., 2024), Gemma 2 (Rivière et al., 2024), and Qwen 3 (Yang et al., 2025). This universality aligns with prior observations of shared structures across models (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025), while revealing a distinct form of representation collapse (Hua et al., 2021; Jing et al., 2022) at the level of attention outputs. We trace the origin of this low-rank structure to the anisotropy of the output projection matrix $W^{O}$, which further compresses multi-head activations into a lower-dimensional subspace.
In Section 5, we investigate how the low-rank structure of attention outputs interacts with SAE training. By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low effective dimensionality strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations. Drawing inspiration from Phan et al. (2025)’s principal component initialization for the first network layer, we propose Active Subspace Initialization, which aligns SAE features with the active subspace of activations, substantially reducing dead features while improving reconstruction. Following Lindsey et al. (2024a) and Gao et al. (2024), we conduct scaling experiments, which further reveal that ASI achieves superior reconstruction across feature counts, and when combined with SparseAdam [^1] See the SparseAdam documentation for details., it achieves the best reconstruction in large scale and reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features trained on Llama-3.1-8B (Dubey et al., 2024).
@@ 2 Related Work
2.1 Representation Collapse in Neural Representations and Low-Rankness in Attention Mechanism
A long line of research has shown that neural representations frequently concentrate in low-dimensional subspaces, forming representation collapse (Hua et al., 2021; Tian et al., 2021; Jing et al., 2022). These works generally focus on visual models trained using self-supervised methods.
Within attention mechanisms, prior work has investigated various notions of “low-rankness”: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach and Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
@@ 2.1 Representation Collapse in Neural Representations and Low-Rankness in Attention Mechanism
2.2 Superposition Hypothesis and Sparse Dictionary Learning Methods
The superposition hypothesis posits that neurons encode multiple non-orthogonal underlying features (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024). Motivated by this view, a variety of sparse dictionary learning methods have been developed for interpretability, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025). These approaches decompose activations into sparse combinations of learned features while differing in their mechanisms for predicting or approximating feature activations. Their successful application across a wide range of model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024) highlights their practical effectiveness for interpretability; however, they do not constitute direct hypothesis tests of the superposition hypothesis, which remain active topics of debate (Sharkey et al., 2025).
@@ 2.4 PCA-Inspired Network Initialization
A common practice applies PCA to input data for dimensionality reduction before network training (Hastie et al., 2009; Montavon et al., 2012; Jolliffe, 1986; Bishop and Nasrabadi, 2007). Recently, Phan et al. (2025) proposed PCsInit, which initializes the first layer weights of networks with top principal components of data—embedding the PCA transform directly into the network. This provides the model with a superior parameter set (Gu et al., 2025), boosting performance by construction.
Figure 2: Across layers, model families and datasets, attention outputs exhibit dramatically lower effective rank than residual streams and MLP outputs, indicating that the attention layer writing into a low dimensional subspace of residual stream is a universal phenomenon. Details in Section 4.1. (left) Evaluation of Llama-3.1-8B on SlimPajama (Soboleva et al., 2023) dataset. (mid) Middle-layer analysis across model families on SlimPajama dataset. (right) Middle-layer analysis of Llama-3.1-8B across datasets.
@@ 3.1 Multi-Head Self-Attention and Notations
$Q_{i}=XW^{Q}_{i}$ $,\qquad K_{i}=XW^{K}_{i},\qquad V_{i}=XW^{V}_{i},$
$W^{Q}_{i}$ $,\,W^{K}_{i},\,W^{V}_{i}\in\mathbb{R}^{d\times d_{h}},$
@@
$\small A_{i}=\mathrm{softmax}\left(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{h}}}\right),\quad Z_{i}=A_{i}V_{i}\in\mathbb{R}^{n\times d_{h}}.$
Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated outputs of all attention heads (Nanda and Bloom, 2022). The final attention output is obtained by applying the output projection:
$O$ $=ZW^{O}=[Z_{1},\dots,Z_{H}]\,[W^{O}_{1},\dots,W^{O}_{H}]^{\top}$
$=\sum_{i=1}^{H}Z_{i}W^{O}_{i}=\sum_{i=1}^{H}O_{i},$
where each $W^{O}_{i}\in\mathbb{R}^{d_{h}\times d}$ is the corresponding submatrix of $W^{O}\in\mathbb{R}^{d\times d}$ associated with each head $i$.
This formulation makes explicit that $O$ is the sum of the outputs from all heads, where each head produces a rank-$d_{h}$ output that is projected into the residual stream space through its corresponding $W^{O}_{h}$. Thus, $O$ represents the attention block’s total contribution to the residual stream.
@@ 3.2 TopK Sparse Autoencoders
$\small\hat{x}=W_{d}z+b_{d}.$
The model is trained to minimize the reconstruction loss, optionally augmented with an auxiliary loss to prevent dead latents:
$\small\mathcal{L}_{\text{TopK-SAE}}=\|x-\hat{x}\|_{2}^{2}+\alpha\cdot\mathcal{L}_{\text{aux}},$
@@ 4 Low-Rank Structure of Attention Outputs
We begin by presenting our central empirical finding: in Transformer models, attention outputs consistently display the strongest low-rank structure compared to MLP outputs and residual streams. As shown in Figure 2, attention outputs have a significantly lower effective rank. This phenomenon is remarkably robust, holding across different layers, model families and datasets. Further details regarding the activation sources are provided in Section 4.2 and Appendix A. These observations highlight that the attention block modifies a subspace of the residual stream, while the MLP operates nearly on the full space.
4.1 Quantifying Low-Rankness with Effective Rank
@@
We consider the activation matrix $A\in\mathbb{R}^{n\times d}$, where each row corresponds to the activation vector of a single token. Here, $n$ denotes the number of data points and $d$ the dimensionality of the activation space (e.g., the model’s hidden size). Unless otherwise specified, $\widetilde{A}$ represents mean-centered activations.
To quantify the effective dimensionality of the activations, we adopt the effective rank metric introduced by Roy and Vetterli (2007).
Definition 4.1 (Effective Rank, Roy and Vetterli (2007)).
Let $\widetilde{A}$ be a nonzero matrix with singular value decomposition $\widetilde{A}=U\Sigma V^{\top}$, where $\Sigma=\mathrm{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{r})$ contains singular values in descending order. Define the normalized singular value distribution as
$\small p_{k}=\frac{\sigma_{k}}{\sum_{j=1}^{r}\sigma_{j}},\quad k=1,2,\ldots,r.$
The (Shannon) entropy of this distribution is
$\small H(p_{1},p_{2},\ldots,p_{r})=-\sum_{k=1}^{r}p_{k}\log p_{k}.$
Then, the effective rank of $\widetilde{A}$ is defined as
$\small\mathrm{erank}(\widetilde{A})=\exp\!\big\{H(p_{1},p_{2},\ldots,p_{r})\big\}.$
Intuitively, the effective rank captures how evenly the singular values are distributed—higher values indicate a more isotropic spectrum, whereas lower values reflect concentration along a few dominant directions. Fraction of effective rank used in Figure 2 means effective rank divided by the dimension of activation space.
Following Bricken et al. (2023b) and Rajamanoharan et al. (2024a), we compute the fraction of downstream loss recovered by varying the number of retained components. We decompose the activations into singular value components and take the language model loss under full activation ablation as a baseline. As components are progressively reintroduced, we report the fraction of this ablated loss that is recovered. See Appendix B for a formal definition.
These quantitative measures complement our core analyses by providing a numerical characterization of the low-rank structure present in activations.
4.2 Experiment Settings
For each activation type, we collect a total of 10 million activation vectors. [^3] In extremely rare cases, outlier activations inflate variance along certain directions, potentially biasing variance-based dimensionality estimates. To mitigate this, we exclude activations whose norms exceed $5\sigma$ from the mean. We empirically verified this sample size suffices to ensure stable and reproducible singular spectrum analyses in Appendix C.
Unless otherwise specified, all experiments in Section 4 run on the middle layer of Llama-3.1-8B (layer 15, zero-indexed), using SlimPajama dataset.
4.3 Empirical Evidence of Low-Rank Structure
(a) Singular Value Spectra.
(b) Fraction of Loss Recovered.
Figure 3: (a) The attention output is the most low-rank, as indicated by the sharpest decay in singular values. (b) Fraction of loss recovered using varying numbers of top singular components.
We draw our findings from three lines of evidence:
Low Effective Rank of Attention Output
Attention outputs have a effective rank of around 60% of the total dimensionality. In contrast, MLP outputs and the residual streams show much higher effective rank around 90% (Figure 2).
Rapid Singular Spectral Decay in Attention Output
This is quantitatively evidenced by the number of components retaining significant energy: only 74.7% singular values exceed 1% [^4] The resolution of bfloat16 (Kalamkar et al., 2019) is 0.01. of the maximum in attention output, versus 100.0% for MLP output and residual stream (Figure 3(a)).
Efficient Downstream Loss Recovery
Compared to zero ablation, attention output requires only 39.1% of dimensions to recover over 99% of the downstream loss, versus 95.3% and 96.9% of the dimensions for MLP outputs and residual streams to recover the same proportion (Figure 3(b)).
More results of these metrics across different layers, models, datasets and activation positions are shown in Appendix D.
4.4 The Output Projection Matrix Further Reduces the Effective Rank of Attention Outputs
Among all activation types, attention outputs consistently exhibits the most rapid singular spectral decay. To investigate whether this low-rank structure originates from the attention heads outputs ($Z$), the output projection matrix ($W^{O}$), or their interaction, we perform a decomposition-based analysis.
Recall that the attention output is computed as $O=ZW^{O}$, where $Z\in\mathbb{R}^{n\times d}$ is the concatenated output of all attention heads, and $W^{O}\in\mathbb{R}^{d\times d}$ is a learned linear projection. To understand how the singular value spectra of $O$ is shaped, we analyze the variance [^5] For zero-mean activations, singular values correspond to the standard deviations of activations along the associated singular directions. of $O$ along a unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
@@ 4.4 The Output Projection Matrix Further Reduces the Effective Rank of Attention Outputs
This expression highlights that the variance along $e$ is determined by two factors: the norm of $W^{O}\hat{e}$ and the variance of $Z$ along $W^{O}\hat{e}$. Specifically, we can rewrite the variance as:
$\small\mathrm{Var}(O\hat{e})=\mathrm{Var}(Z\hat{v})\cdot\|v\|_{2}^{2}\;\text{where}\;v=W^{O}\hat{e},\ \hat{v}=\frac{v}{\|v\|_{2}}.$
@@
We compute and visualize the singular values of attention output $O$ and these two contributions, as shown in Figure 4. This analysis reveals that the low-rank structure of attention outputs is strongly influenced by $W^{O}$, which further compresses $Z$ into a lower-dimensional subspace. The analysis of the effective rank of $Z$ in Figure 11 further supports this conclusion. From a mechanistic perspective, an intuitive explanation is that although each attention head contributes a $d_{\text{head}}$-dimensional subspace, the superposition of attention heads (Jermyn et al., 2024; He et al., 2025) inevitably induces overlaps among these subspaces. Let $O_{i}$ denote the output of the $i^{\text{th}}$ attention head. Consequently, the dimension of the MHSA output satisfies
$\dim\!\left(\bigcup_{i}\text{span}(O_{i})\right)$ $\;\leq\;\sum_{i}\dim\!\big(\text{span}(O_{i})\big)$
$\;=\;d_{\text{head}}\cdot n_{\text{head}}$
$\;=\;d_{\text{model}}\qquad\text{(in standard MHSA)}.$
@@
Figure 5: The number of dead features (left) and the effective rank (mid) of each activation in Llama-3.1-8B, shows a surprising consistency (right): activations with lower effective rank have more dead features, corresponding to all layers of attention output and last two layers of MLP output.
Figure 6: After using ASI, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different $d_{init}$ for activations with a full space dimension of 4096. All experiments repeat 3 times using different random seeds and error bars indicate mean ± std.
Figure 7: At a fixed number of features ($n=32768$), ASI (TopK SAE with Active Subspace Init) achieves a better reconstruction-sparsity trade-off than TopK (standard TopK SAE) and AuxK (TopK SAE with auxiliary loss). A similar trend is observed in its impact on Delta LM Loss. All experiments repeat 3 times using different random seeds and show the mean. The results of std are in Appendix G due to resolution constraints.